Choose question
1.a ball of mass 1kg is thrown at 8m/s and kicked a building.if it turns back at 6m/s,calculate the magnitude of change in momentum.
A.2Kg m/s B.14Kg m/s C.4Kg m/d D.0K m/s
2.if a car is moving uniformly at 8m/s,then which of the following changes uniformly?
A.acceleration B.velocity C.A&B D.none
3.car A traveling at velocity of 30m/s relative to the earth is overtaken by another car B moving at 40m/s in the same direction.what is the velocity of car B relative to car A?
A.10m/s B.70m/s C.59m/s D.20m/s
4.an automobile accelerated for 5sec.and attain a maximum speed of 10m/s.it continued at this speed 4sec.if the automobile decelerates for the next 6sec before it stops,calculate the total distance covered.
A.65m B.70m C.95m D.100m
5.calculate the force required to accelerate a 2Kg object at 5m/s2.
A.10N B.2.5N C.3N D.7N
6.determine the weight of 50 bricks;each brick has a mass of 2.5Kg
A.1200N B.2000N C.1250N D.4000N
1
Expert's answer
2020-12-21T11:02:49-0500

1.


"\Delta p=1(6+8)=14\frac{kgm}{s}"

Answer: B.14Kg m/s


2. Answer: D.none


3.


"\Delta v=40-30=10\frac{m}{s}"

Answer: A.10m/s

4.


"s=0.5(10)(5)+(10)(4)+0.5(10)(6)=95\ m"

Answer: C.95m 

5.


"F=ma=(2)(5)=10\ N"

Answer: A.10N 

6.


"W=nmg=(50)(2.5)(10)=1250\ N"

Answer: C.1250N